# Phase error rate estimation with basis-efficiency mismatch for decoy-state BB84

### v1

@@ B.1 Random Sampling
Serfling basically states that if one chooses a random set of positions, then the fraction of $1$s in those positions gives us a good estimate of the fraction of $1$s in the remaining positions. However, observe that the sampling procedure in the protocol from Sections IV.3 and V.4 does not actually choose a random subset of fixed-length for testing. Instead, the protocol decides to map each conclusive round to test or key in an IID manner. Therefore, the application of the Serfling bound is not straightforward. In the following lemma, we show how the Serfling bound can still be rigorously used.
@@
B.2 Sampling with imperfect detectors
We now turn our attention to proving Lemma 3, which is the main statement utilized in extending the EUR approach to imperfect detectors in Section V. We start by proving Lemmas 8 and 4 which we use later in the proof of Lemma 9.
@@ B.2 Sampling with imperfect detectors
The proof follows after noting that Eqs. 92 and 93 $\implies$ Eq. 91. ∎
@@
The proof follows from the observation that Eqs. 94, 95 and 96 $\implies$ Eq. 34. ∎
B.3 Sampling with independent imperfect detectors
@@
Appendix C Combining bounds
In this section, we will combine Eqs. 30, 33, 38 and 41 and obtain Eq. 42. This process is simply some cumbersome algebra and the use of the union bound for probabilities.
Combining Eqs. 30 and 33, we obtain
@@ Appendix C Combining bounds
Combining Eqs. 99 and 38 by setting $f=\bm{e^{\text{obs}}_{X}}+\gamma^{\varepsilon_{\text{AT-a}}}_{\text{serf}}(\tilde{n}_{X},\tilde{n}_{K})$ in Eq. 38, we get
@@
$\leq\Pr(\bm{\tilde{e}^{\text{key}}_{XX}}\geq\bm{e^{\text{obs}}_{X}}+\gamma^{\varepsilon_{\text{AT-a}}}_{\text{serf}}(\tilde{n}_{X},\tilde{n}_{K}))_{|\Omega_{(\tilde{n}_{X},\tilde{n}_{K})}}+\frac{\varepsilon_{\text{AT-b}}^{2}}{2}$
@@
Combing Eqs. 100 and 41, we obtain
@@
Appendix D Decoy Analysis
@@
In the decoy-state QKD protocol of Section VI, fix an outcome $O$, and intensity $\mu_{k}$. Then, we have
@@
Appendix E Detector Model Calculations
@@
In this section we will compute upper bounds on the $\delta_{1},\delta_{2}$. To do so, we follow the procedure from Section VII.2.
E.1 Computing POVMs
We first calculate the different POVM elements from Section VII.2 assuming that Alice and Bob pick the same basis $b$,
@@ E.1 Computing POVMs
$\tilde{F}_{(b,b),(\text{con})}=$ $\mathrm{I}_{A}\otimes\sum_{N_{0},N_{1}=0}^{\infty}(1-(1-d_{b_{0}})(1-d_{b_{1}})(1-\eta_{b_{0}})^{N_{0}}(1-\eta_{b_{1}})^{N_{1}})\outerproduct{N_{0},N_{1}}{N_{0},N_{1}}_{b}$ (108)
$\tilde{F}_{(b,b),(\bot)}=$ $\mathrm{I}_{A}\otimes\sum_{N_{0},N_{1}=0}^{\infty}(1-d_{b_{0}})(1-d_{b_{1}})(1-\eta_{b_{0}})^{N_{0}}(1-\eta_{b_{1}})^{N_{1}}\outerproduct{N_{0},N_{1}}{N_{0},N_{1}}_{b}$
@@
We can choose
It is straightforward to verify that $\tilde{F}$ satisfies our requirement $\tilde{F}\geq\tilde{F}_{(b_{A},b_{B}),(\text{con})}$ for all choices of $b_{A},b_{B}$. Then, we calculate the POVM elements from Eq. 76 as
${F}_{(b),(\bot)}$ $=\mathrm{I}_{A}\otimes\sum_{N_{0}+N_{1}\neq 0}^{\infty}\left(1-\frac{1-(1-d_{b_{0}})(1-d_{b_{1}})(1-\eta_{b_{0}})^{N_{0}}(1-\eta_{b_{1}})^{N_{1}}}{(1-(1-d_{\text{max}})^{2}(1-\eta_{\text{max}})^{N_{0}+N_{1}})}\right)\outerproduct{N_{0},N_{1}}{N_{0},N_{1}}_{b}.$
E.2 Computing $\delta_{1}$
@@
We write
@@
where Eq. 116 is a consequence of the triangle inequality. Let us focus on a fixed value of $N=N_{1}+N_{2}$. In this case, combining Eqs. 113, 114 and 116 allows us to obtain
$\delta_{1}^{(N)}$ $\leq 4\sum_{b=X,Z}\Biggl{|}\sqrt{\frac{1-(1-(1-d_{b_{0}})(1-d_{b_{1}})(1-\eta_{b_{0}})^{N_{0}}(1-\eta_{b_{1}})^{N_{1}}}{1-(1-\eta_{\text{max}})^{N_{0}+N_{1}}}}$ (117)
$-\frac{\sqrt{1-(1-d_{\text{max}})^{2}(1-\eta_{\text{max}})^{N_{0}+N_{1}}}+\sqrt{1-(1-d_{\text{min}})^{2}(1-\eta_{\text{min}})^{N_{0}+N_{1}}}}{2\sqrt{1-(1-\eta_{\text{max}})^{N_{0}+N_{1}}}}\Biggl{|}$
@@
where the final inequality above follows from the monotonicity of the expression inside the modulus with respect to $\eta_{b_{0}},\eta_{b_{1}},d_{b_{0}},d_{b_{1}}$. However, as described in Ref. [52, Section III C] we can renormalize detection efficiencies and treat the common loss in detectors as part of the channel. Thus, without loss of generality, we consider one of the detectors to be lossless. This can be thought of as setting $\eta_{\text{max}}\rightarrow 1$ and $\eta_{\text{min}}\rightarrow\eta_{\text{min}}/\eta_{\text{max}}$. This significantly simplifies the calculations. Combining Eqs. 112, 114, 116, 117 and 111, we obtain
$\delta_{1}$ $\leq 4\max\left\{1-\sqrt{\frac{1-(1-d_{\text{min}})^{2}}{1-(1-d_{\text{max}})^{2}}},\max_{N\geq 1}\absolutevalue{1-\sqrt{1-(1-d_{\text{min}})^{2}(1-r_{\eta})^{N}}}\right\}$ | } (118)
$\leq 4(1-\sqrt{1-(1-d_{\text{min}})^{2}(1-r_{\eta})})\qquad\qquad(\text{when $d_{\text{min}}\neq 0$}),$
@@
E.3 Computing $\delta_{2}$
We can now also compute $\delta_{2}=\norm{\mathrm{I}-{F}_{(Z),(\text{con})}}_{\infty}$ in a similar way, using Eqs. 110 and 77. Again, we have $\delta_{2}=\max_{N}\delta_{2}^{(N)}$, where

### v2

@@ C.1 Random Sampling
Serfling basically states that if one chooses a random set of positions, then the fraction of $1$s in those positions gives us a good estimate of the fraction of $1$s in the remaining positions. However, observe that the sampling procedure in the protocol from Sections III.3 and IV.4 does not actually choose a random subset of fixed-length for testing. Instead, the protocol decides to map each conclusive round to test or key in an IID manner. Therefore, the application of the Serfling bound is not straightforward. In the following lemma, we show how the Serfling bound can still be rigorously used.
@@
C.2 Sampling with imperfect detectors
We now turn our attention to proving Lemma 3, which is the main statement utilized in extending the EUR approach to imperfect detectors in Section IV. We start by proving Lemmas 8 and 4 which we use later in the proof of Lemma 9.
@@ C.2 Sampling with imperfect detectors
The proof follows after noting that Eqs. 87 and 88 $\implies$ Eq. 86. ∎
@@
The proof follows from the observation that Eqs. 89, 90 and 91 $\implies$ Eq. 24. ∎
C.3 Sampling with independent imperfect detectors
@@
Appendix D Combining bounds
In this section, we will combine Eqs. 20, 23, 28 and 31 and obtain Eq. 32. This process is simply some cumbersome algebra and the use of the union bound for probabilities.
Combining Eqs. 20 and 23, we obtain
@@ Appendix D Combining bounds
We will now combine Eqs. 94 and 28. To do so we will additionally need to condition on $e^{\text{obs}}_{X}$. However, note that Eq. 28 remains true with this additional conditioning (because $e^{\text{obs}}_{X}$ is observed on a different set of rounds). Thus we obtain
@@
$=\sum_{e^{\text{obs}}_{X}}\Pr(\Omega_{(e^{\text{obs}}_{X})}|\Omega_{(\tilde{n}_{X},\tilde{n}_{K})})\Pr(\bm{\tilde{e}^{\text{key}}_{ZX}}\geq e^{\text{obs}}_{X}+\gamma^{\varepsilon_{\text{AT-a}}}_{\text{serf}}(\tilde{n}_{X},\tilde{n}_{K})+\delta_{1}+\gamma^{\varepsilon_{\text{AT-b}}}_{\text{bin}}(\tilde{n}_{K},\delta_{1}))_{|\Omega_{(\tilde{n}_{X},\tilde{n}_{K},e^{\text{obs}}_{X})}}$
$\leq\sum_{e^{\text{obs}}_{X}}\Pr(\Omega_{(e^{\text{obs}}_{X})}|\Omega_{(\tilde{n}_{X},\tilde{n}_{K})})\Pr(\bm{\tilde{e}^{\text{key}}_{XX}}\geq e^{\text{obs}}_{X}+\gamma^{\varepsilon_{\text{AT-a}}}_{\text{serf}}(\tilde{n}_{X},\tilde{n}_{K}))_{|\Omega_{(\tilde{n}_{X},\tilde{n}_{K},e^{\text{obs}}_{X})}}+\frac{\varepsilon_{\text{AT-b}}^{2}}{2}$
$=\Pr(\bm{\tilde{e}^{\text{key}}_{XX}}\geq\bm{e^{\text{obs}}_{X}}+\gamma^{\varepsilon_{\text{AT-a}}}_{\text{serf}}(\tilde{n}_{X},\tilde{n}_{K}))_{|\Omega_{(\tilde{n}_{X},\tilde{n}_{K})}}+\frac{\varepsilon_{\text{AT-b}}^{2}}{2}$
@@
where the first equality follows from the definition of conditional probability. The second inequality is obtained by setting $e=e^{\text{obs}}_{X}+\gamma^{\varepsilon_{\text{AT-a}}}_{\text{serf}}(\tilde{n}_{X},\tilde{n}_{K})$ in Eq. 28, the third equality follows from the definition of probability and the final inequality follows from Eq. 94
Combing Eqs. 95 and 31, we obtain
@@
Appendix E Decoy Analysis
@@
In the decoy-state QKD protocol of Section V, fix an outcome $O$, and intensity $\mu_{k}$. Then, we have
@@
Appendix F Variable-length security proof for decoy-state BB84
In this appendix, we will prove the following theorem regarding the variable-length security of the decoy-state BB84 protocol.
See 3
Proof.
Again, the proof is similar to that of Theorem 1 with some differences. The first part of the proof is identical to that of Theorem 1 and we do not repeat it here. Thus, this proof consists of two parts. In the first part, we use Eq. 38 along with the entropic uncertainty relations to bound the smooth min-entropy of the raw key register. Here the main difference is the use of entropic chain rules to isolate the single-photon component of the raw key, before applying the EUR statement. In the second part, we use the obtained bound to prove the variable-length security statement. Here the main difference is due to to the presence of the $n_{K,{1}}$ in the event $\Omega_{(n_{X,\mu_{\vec{k}}},n_{K,\mu_{\vec{k}}},e^{\text{obs}}_{X,\mu_{\vec{k}}},e^{\text{obs}}_{Z},n_{K,{0}},n_{K,{1}})}=\widetilde{\Omega}$, which is handled differently. Let us first focus on the event $\Omega_{(n_{X,\mu_{\vec{k}}},n_{K,\mu_{\vec{k}}},e^{\text{obs}}_{X,\mu_{\vec{k}}},e^{\text{obs}}_{Z},n_{K,{0}},n_{K,{1}})}=\widetilde{\Omega}$. (In the remainder of this proof, we identify $\widetilde{\Omega}=\Omega_{(n_{X,\mu_{\vec{k}}},n_{K,\mu_{\vec{k}}},e^{\text{obs}}_{X,\mu_{\vec{k}}},e^{\text{obs}}_{Z},n_{K,{1}})}$ for brevity.)
F.0.1 Bounding the smooth min entropy
Similar to the proof of Theorem 1, we will use $\kappa(\widetilde{\Omega})$, to denote the probability of our computed bounds failing conditioned on the event $\widetilde{\Omega}$. We will see that the events that we will need to condition on are given by $\Omega_{(n_{X,\mu_{\vec{k}}},n_{K,\mu_{\vec{k}}},e^{\text{obs}}_{X,\mu_{\vec{k}}},e^{\text{obs}}_{Z},n_{K,{1}})}$. Even though we do not have access to the value $n_{K,{1}}$, we will see that this is the “right” event to condition on. We do not generate key from the zero-photon component $n_{K,{0}}$, since it has little impact on key rates for typical use cases. To see how one may do so, see Remark 19. As in Appendix B, we define
which when combined with Eq. 38 implies
Thus, the average probability of either of our bounds failing (for any attack undertaken by Eve) is small.
where we used [54, Lemma 6.7] to isolate the single-photon contribution to the key in the first inequality, and the entropic uncertainty relations [1] followed by the same series of steps as in the proof of Theorem 1 to replace the smooth max entropy term with the bound on the phase error rate. This is the required bound on the smooth min entropy of the raw key.
Remark 19.
$=\Delta_{1}+\Delta_{2}$
Here, we used the leftover-hashing lemma [2, Proposition 9] on the sub-normalized state $(\rho_{|\widetilde{\Omega}})_{\wedge\Omega_{\text{EV}}}$ for the first inequality, and [2, Lemma 10] to get rid of the sub-normalized conditioning ($\wedge\Omega_{\text{EV}}$) in the smooth min entropy term in the second inequality. We used [54, Lemma 6.8] to split off the error-correction information ($\lambda_{\text{EC}}(n_{X,\mu_{\vec{k}}},n_{K,\mu_{\vec{k}}},e^{\text{obs}}_{X,\mu_{\vec{k}}},e^{\text{obs}}_{Z})$) and error-verification information ($\log(2/\varepsilon_{\text{EV}})$) in the third inequality. We used the bound on the smooth min entropy from Eq. 105 for the fourth inequality. The fifth equality follows by replacing the values of $l(n_{X},n_{K},e^{\text{obs}}_{X},e^{\text{obs}}_{Z})$ and $\lambda_{\text{EC}}(n_{X},n_{K},e^{\text{obs}}_{X},e^{\text{obs}}_{Z})$ from Eq. 53. We use the concavity of the square root function and Jensen’s inequality for the sixth inequality to pull the sum over the events and the probability inside the square root, while the seventh equality follows simply from the definition of conditional probabilities and $\kappa$ (Eq. 103).
Combining our bounds on $\Delta_{1}$ and $\Delta_{2}$ from Eqs. 109 and 108, we obtain
$\leq\varepsilon_{\text{PA}}+2\varepsilon_{\text{AT}}.$
Notice that our proofs of variable-length security (Theorems 1 and 3) rely on obtaining a suitable bound on the smooth min entropy of the raw key register, with a suitable smoothing parameter, for suitable events. In particular, the smoothing parameter averaged over all events satisfies certain bounds. However, the theorem statements so far have been specific to BB84 and decoy-state BB84. We state the following technical theorem regarding variable-length security which is applicable to generic protocols, as long as suitable bounds on the smooth min entropy can be obtained.
Theorem 4.
In a QKD protocol, let $\Omega_{(i,j)}$ denote well-defined events (we use $i$ for observed events and $j$ for unobserved events) that can take place. Let $\vec{Z}$ denote the raw key register, let $\vec{E}$ denote Eve’s quantum system, and let $\vec{C}$ denote public announcements (excluding error-correction and error-verification). Let the protocol be such that it produces a key of length $l(i)$ bits, using $\lambda_{\text{EC}}(i)$ bits for error-correction and $\log(2/\varepsilon_{\text{EV}})$ bits for error-verification, upon the observed event $\Omega_{(i)}\wedge\Omega_{\text{EV}}$. For each $i$, let $S_{i}$ denote a subset of possible values of $j$, and let $\kappa_{(i,j)}\geq 0$ be a set of values such that
$H^{\sqrt{\kappa_{(i,j)}}}_{\text{min}}(\vec{Z}|\vec{E}\vec{C})_{\rho|\Omega_{(i,j)}}\geq\beta_{i}\qquad\qquad\forall i,\forall j\in S_{i}$ (111)
$\sum_{i}\sum_{j\in S_{i}}\Pr(\Omega_{(i,j)})\kappa_{(i,j)}+\sum_{i}\sum_{j\notin S_{i}}\Pr(\Omega_{(i,j)})\leq\varepsilon_{\text{AT}}^{2}$
$l(i)\coloneqq\max\big{\{}\beta_{i}-\lambda_{\text{EC}}(i)-2\log(1/2\varepsilon_{\text{PA}})-\log(2/\varepsilon_{\text{PA}}),0\big{\}}.$
Then the QKD protocol is $(2\varepsilon_{\text{AT}}+\varepsilon_{\text{PA}}+\varepsilon_{\text{EV}})$-secure.
Proof Sketch. The proof follows an identical series of steps as the proof of Theorem 3, and we do not repeat it here. This can be seen by identifying $i$ with observed events in the protocol (analogous to $(n_{X,\mu_{\vec{k}}},n_{K,\mu_{\vec{k}}},e^{\text{obs}}_{X,\mu_{\vec{k}}},e^{\text{obs}}_{Z})$), and $j$ with events that are not directly observed in the protocol (analogous to $n_{K,{1}}$). One identifies the set $S_{i}$ with values of $j$ that satisfy certain bounds with high probability.
Intuitively the conditions in Eq. 111 split all possible combinations of $(i,j)$ into two sets, depending on whether $j\in S_{i}$ or $j\notin S_{i}$. If $j\in S_{i}$, then a suitable bound $\beta_{i}$ on the min entropy is known. This bound is utilized in the leftover hash lemma, and the trace distance for QKD security is bounded in the same manner as Eq. 109. The bound obtained in this case is given by $\varepsilon_{\text{PA}}+2\sqrt{\sum_{i,j\in S_{i}}\Pr(\Omega_{(i,j)})\kappa_{(i,j)}}$. If $j\notin S_{i}$, then the trace distance for QKD security is bounded in the same manner as Eq. 108. The bound obtained is given by $\sum_{i,j\notin S_{i}}\Pr(\Omega_{(i,j)})$. These two bounds can be combined as in Eq. 110.
Note that if $j$ is set to be a trivial value, then given the suitable bound on the min entropy, we recover Theorem 1.
In general, when dealing with events, one must ensure that all events considered are well-defined, i.e, there exists (in theory) a classical register that determines whether the event occured or did not occur [2, Section 2].
Appendix G Detector Model Calculations
@@
In this section we will compute upper bounds on the $\delta_{1},\delta_{2}$. To do so, we follow the recipe from Section VI.1.
G.1 Computing POVMs
Recall that Alice and Bob’s joint POVM $\{\Gamma_{(b_{A},b_{B}),(\neq)},\Gamma_{(b_{A},b_{B}),(=)},\Gamma_{(b_{A},b_{B}),(\bot)}\}$ is given in Eq. 58. We can choose
It is straightforward to verify that $\tilde{F}$ satisfies our requirement $\tilde{F}\geq\Gamma_{(b_{A},b_{B}),(\neq)}+\Gamma_{(b_{A},b_{B}),(=)}=\mathrm{I}-\Gamma_{(b_{A},b_{B}),(\bot)}$ for all choices of $b_{A},b_{B}$. Then, we calculate the POVM elements from Eq. 55 as
@@ G.1 Computing POVMs
${F}_{(b),(\bot)}$ $=\mathrm{I}_{A}\otimes\sum_{N_{0},N_{1}=0}^{\infty}\left(1-\frac{1-(1-d_{b_{0}})(1-d_{b_{1}})(1-\eta_{b_{0}})^{N_{0}}(1-\eta_{b_{1}})^{N_{1}}}{(1-(1-d_{\text{max}})^{2}(1-\eta_{\text{max}})^{N_{0}+N_{1}})}\right)\outerproduct{N_{0},N_{1}}{N_{0},N_{1}}_{b}$
@@
G.2 Computing $\delta_{1}$
@@
First, we compute $\delta_{1}^{(0)}$ directly as
$\frac{\delta_{1}^{(0)}}{2}=$ $\norm{\mathrm{I}_{A}\otimes\outerproduct{0,0}{0,0}\left(\sqrt{{F}_{(Z),(\text{con})}}G^{\text{con}}_{(X),(\neq)}\sqrt{{F}_{(Z),(\text{con})}}-\sqrt{{F}_{(X),(\text{con})}}G^{\text{con}}_{(X),(\neq)}\sqrt{{F}_{(X),(\text{con})}}\right)\mathrm{I}_{A}\otimes\outerproduct{0,0}{0,0}}_{\infty}$
$=$ |1($\absolutevalue{\frac{1-(1-d_{Z_{0}})(1-d_{Z_{1}})}{1-(1-d_{\text{max}})^{2}}-\frac{1-(1-d_{X_{0}})(1-d_{X_{1}})}{1-(1-d_{\text{max}})^{2}}}\max\left\{\frac{d_{X_{1}}(2-d_{X_{0}})}{2(1-(1-d_{X_{0}})(1-d_{X_{1}}))},\;\frac{d_{X_{0}}(2-d_{X_{1}})}{2(1-(1-d_{X_{0}})(1-d_{X_{1}}))}\right\}$
$=$ |$\absolutevalue{\frac{(1-d_{X_{0}})(1-d_{X_{1}})-(1-d_{Z_{0}})(1-d_{Z_{1}})}{1-(1-d_{\text{max}})^{2}}}\max\left\{\frac{d_{X_{1}}(2-d_{X_{0}})}{2(1-(1-d_{X_{0}})(1-d_{X_{1}}))},\;\frac{d_{X_{0}}(2-d_{X_{1}})}{2(1-(1-d_{X_{0}})(1-d_{X_{1}}))}\right\}$
$\leq$ $\left(1-\frac{1-(1-d_{\text{min}})^{2}}{1-(1-d_{\text{max}})^{2}}\right)\frac{d_{\text{max}}(2-d_{\text{min}})}{2(1-(1-d_{\text{min}})^{2})}.$ (115)
To bound $\delta_{1}^{(N)}$ for $N\neq 0$, we write
@@
where Eq. 120 is a consequence of the triangle inequality. Let us focus on a fixed value of $N=N_{0}+N_{1}\neq 0$. In this case, combining Eqs. 117, 118 and 120 allows us to obtain
$\delta_{1}^{(N)}$ $\leq 4\sum_{b=X,Z}\Biggl{|}\sqrt{\frac{1-(1-d_{b_{0}})(1-d_{b_{1}})(1-\eta_{b_{0}})^{N_{0}}(1-\eta_{b_{1}})^{N_{1}}}{1-(1-d_{\text{max}})^{2}(1-\eta_{\text{max}})^{N_{0}+N_{1}}}}$ (121)
$-\frac{\sqrt{1-(1-d_{\text{max}})^{2}(1-\eta_{\text{max}})^{N_{0}+N_{1}}}+\sqrt{1-(1-d_{\text{min}})^{2}(1-\eta_{\text{min}})^{N_{0}+N_{1}}}}{2\sqrt{1-(1-d_{\text{max}})^{2}(1-\eta_{\text{max}})^{N_{0}+N_{1}}}}\Biggl{|}$
@@
where the final inequality above follows from the monotonicity of the expression inside the modulus with respect to $\eta_{b_{0}},\eta_{b_{1}},d_{b_{0}},d_{b_{1}}$. However, as described in Ref. [52, Section III C] we can renormalize detection efficiencies and treat the common loss in detectors as part of the channel. Thus, without loss of generality, we consider one of the detectors to be lossless. This can be thought of as setting $\eta_{\text{max}}\rightarrow 1$ and $\eta_{\text{min}}\rightarrow\eta_{\text{min}}/\eta_{\text{max}}$. This significantly simplifies the calculations. Combining Eqs. 116, 118, 120, 121 and 114, we obtain
$\delta_{1}^{(N)}$ $\leq 4\absolutevalue{1-\sqrt{1-(1-d_{\text{min}})^{2}(1-r_{\eta})^{N}}},\qquad(N\geq 1)$ (122)
Combining Eqs. 122 and 115, we obtain
$\delta_{1}$ $\leq\max\left\{\left(1-\frac{1-(1-d_{\text{min}})^{2}}{1-(1-d_{\text{max}})^{2}}\right)\frac{d_{\text{max}}(2-d_{\text{min}})}{1-(1-d_{\text{min}})^{2}},\absolutevalue{1-\sqrt{1-(1-d_{\text{min}})^{2}(1-r_{\eta})}}\right\},$ (123)
@@
G.3 Computing $\delta_{2}$
We can now also compute $\delta_{2}=\norm{\mathrm{I}-{F}_{(Z),(\text{con})}}_{\infty}$ in a similar way, using Eqs. 113 and 56. Again, we have $\delta_{2}=\max_{N}\delta_{2}^{(N)}$, where